# Who Said Neural Networks Aren’t Linear?

### v1

@@ Abstract
[…]
@@ 1 Introduction
In contrast, non-linear systems, while more expressive, often defy such a structure. Iterating nonlinear mappings can quickly lead to intractable dynamics; inversion may be ill-posed or undefined; and even simple compositional questions lack closed-form answers (Strogatz, 2024). Neural networks, the dominant modeling tool in modern machine learning, are famously nonlinear, placing their analysis and manipulation outside the reach of classical linear algebra (Hornik et al., 1989). As a result, tasks that are trivial in the linear setting, such as projecting onto a subspace, enforcing idempotency, or collapsing iterative procedures, become major challenges in the nonlinear regime that require engineered loss functions and optimization schemes.
This raises a natural and fundamental question: can we reinterpret conventionally nonlinear neural networks as linear operators in disguise? If so, we would gain access to the full arsenal of linear methods without sacrificing the expressive power of nonlinear function classes.
@@ 2 Linearizer Framework
Linearity is not an absolute concept; it is a property defined relative to a pair of vector spaces, $f:\mathcal{X}\to\mathcal{Y}$. This begs the question: is it possible to identify a pair of spaces $\mathcal{X}$ and $\mathcal{Y}$, for which a function that is nonlinear over standard Euclidean space is, in fact, perfectly linear? To answer this, we must first deconstruct what defines a vector space. It is a structure comprised of a set of vectors (e.g. $\mathbb{R}^{N}$), a field of scalars (e.g., $\mathbb{R}$), and two fundamental operations: vector addition (++) and scalar multiplication ($\cdot$). We propose that while keeping the set of vectors and the field unchanged, we can redefine the operations in order to induce new vector spaces.
@@ 2.1 Definitions
We introduce a formalism we term as Linearizer, in which the relevant vector spaces are immediately identifiable by construction. The architecture we propose can be trained from scratch or by distillation from an existing model. Our approach is made practical by invertible neural networks Rezende & Mohamed (2015); Dinh et al. (2015; 2017). We build our model by wrapping a linear operator, a matrix $A$, between two such invertible networks, $g_{x}$ and $g_{y}$:
Definition 1 (Linearizer).
@@
Definition 2 (Induced Vector Space Operations).
Let $g:V\to V$ be an invertible function. We define a new set of operations, $\oplus$ and $\odot$, for any two vectors $v_{1},v_{2}\in V$ and any scalar $a\in\mathbb{R}$:
@@ 2.2 Linearity
The input space $\mathcal{X}$ is defined by operations $(\oplus_{x},\odot_{x})$ induced by $g_{x}$, and the output space $\mathcal{Y}$ by $(\oplus_{y},\odot_{y})$ induced by $g_{y}$.
Lemma 1.
$(V,\oplus_{g},\odot_{g})$ is a vector space over $\mathbb{R}$. Proof. See Appendix. A.
Lemma 2.
@@
To prove linearity, we show that $f$ satisfies the superposition principle.
$f(a_{1}\odot_{x}x_{1}\oplus_{x}a_{2}\odot_{x}x_{2})$ $=g_{y}^{-1}(Ag_{x}(g_{x}^{-1}(a_{1}g_{x}(x_{1})+a_{2}g_{x}(x_{2}))))$
@@
$=a_{1}\odot_{y}f(x_{1})\oplus_{y}a_{2}\odot_{y}f(x_{2})\qed$
Figure 1: Left. The Linearizer structure (top) is a linear operation sandwiched between two invertible functions (bottom) Right. Vector addition and scalar multiplication define induced vector spaces for which $f$ is linear.
@@ 2.3 Intuition
The Linearizer can be understood through several intuitive lenses.
Linear algebra analogy. The Linearizer is analogous to an eigendecomposition or SVD. Just as a matrix is diagonal (i.e. simplest) in its eigenbasis, our function $f$ is a simple matrix multiplication in the ”linear basis” defined by $g_{x}$ and $g_{y}$. This provides a natural coordinate system in which to analyze and manipulate the function. For example, repeated application of a Linearizer with a shared basis ($g_{x}=g_{y}=g$) is equivalent to taking a power of its core matrix:
$\mathbb{L}_{\{g,g,A\}}^{\circ N}(x)=\underbrace{f(f(\cdots f}_{N\text{ times}}(x)))=g^{-1}(A^{N}\cdot g(x))=\mathbb{L}_{\{g,g,A^{N}\}}(x)$ (4)
Geometric interpretation. $g_{x}$ can be seen as a diffeomorphism. Equip the input with the pullback metric $g_{x}^{\ast}\langle\cdot,\cdot\rangle_{\mathbb{R}^{N}}$ induced by the Euclidean metric in latent space. With respect to this metric, straight lines in latent correspond to geodesics in input, and linear interpolation in the induced coordinates maps to smooth and semantically coherent curves in input space.
@@ 2.4 Properties
We review how basic properties of linear transforms are expressed by Linearizers.
@@ 2.4.1 Composition
$\text{{Proof. }}(f_{2}\circ f_{1})(x)=g_{z}^{-1}(A_{2}g_{y}(g_{y}^{-1}(A_{1}g_{x}(x))))=g_{z}^{-1}((A_{2}A_{1})\cdot g_{x}(x))\text{\qed}$ (5)
@@ 2.4.2 Inner Product and Hilbert Space
Definition 3 (Induced Inner Product).
@@
Lemma 4 (Induced spaces are Hilbert).
$\langle u,v\rangle_{g}\;:=\;\langle g(u),\,g(v)\rangle_{\mathbb{R}^{n}}.$ (7)
Then $(V,\langle\cdot,\cdot\rangle_{g})$ is a Hilbert space. Proof. See Appendix C
@@ 2.4.3 Transpose
$=\langle x,g_{x}^{-1}(A^{\top}g_{y}(y))\rangle_{g_{x}}=\langle x,f^{\top}(y)\rangle_{g_{x}}.\text{\qed}$ (10)
@@ 2.4.4 Singular Value Decomposition
Proof.
See Appendix B. ∎
@@ 3.1 One-Step Flow Matching
[…]
Figure 2: Comparison between multi-step and one-step flow matching. Panel labels: (a) multi-step Linear FM. (b) one-step Linear flow matching (FM);
[…]
Figure 3: One-step inversion examples: Left and right (in red): original (not generated) data $x_{1}$ and $x_{2}$. Intermediate images obtained by latent interpolation. See Appendix. E.2 for more and higher-resolution results.
Flow matching (Lipman et al., 2023; Song et al., 2021b) trains a network to predict the velocity field that transports noise samples to data points. Traditionally, this velocity must be integrated over time with many small steps, making generation slow. We train flow matching using our Linearizer as the back bone, with a single vector space ($g_{x}=g_{y}=g$). The key property that helps us achieve one-step generation is the closure of linear operators. Applying the model sequentially with some simple linear operations in between steps is a chain of linear operations. This chain can be collapsed to a single linear operator so that the entire trajectory is realized in a single step.
@@
$\mathcal{L}=\mathbb{E}_{x_{0},x_{1},t}\,\bigl\|\,v\ominus f(x_{t},t)\,\bigr\|^{2}=\mathbb{E}_{x_{0},x_{1},t}\,\bigl\|\,g^{-1}\Bigl(g(x_{1})-g(x_{0})-A_{t}g(x_{t})\Bigr)\,\bigr\|^{2}.$ (14)
@@
Thus, what was originally a long sequence of updates collapses into a single multiplication. In practice, higher-order integration such as Runge-Kutta (Runge, 1895; Kutta, 1901) may be used to calculate $B$ more accurately (see Appendix F). $B$ is calculated only once, after training. Then generation requires only a single feedforward activation of our model on noise, regardless of the number of steps discretizing the ODE.
@@
The operator $A$ is produced by an MLP whose input is $t$. To avoid huge matrices multiplications we build A as a low-rank matrix, by multiplying two rectangular matrices. $g$ does not take $t$ as input. Full implementation details are provided in App. G.
@@
Inversion and interpolation.
A fundamental limitation of flow models is that, in contrast to VAEs (Kingma & Welling, 2014), they lack a natural encoder: they cannot map data back into the prior (noise) space, and thus act only as decoders. As a result, inversion methods for diffusion (Dhariwal & Nichol, 2021; Song et al., 2021a; Huberman-Spiegelglas et al., 2024) have become an active research area. However, these techniques are approximate and often suffer from reconstruction errors, nonuniqueness, or computational overhead. Our framework offers a new bridge between diffusion-based models and encoding ability.
Lemma 7 (Moore–Penrose pseudoinverse of a Linearizer).
$f^{\dagger}(y)=g_{x}^{-1}(A^{\dagger}g_{y}(y))$ (19)
Proof.
We verify the four Penrose equations (Penrose, 1955) in Appendix D. ∎
This property enables the exact encoding of data in the latent space. For example, given two data points $x_{1},x_{2}$, we can encode them via $z^{a}=(1-a)\,f^{\dagger}(x_{1})\;+\;a\,f^{\dagger}(x_{2}),$ and decode back by $\hat{x}^{a}=f(z^{a}).$ Fig 3 shows latent interpolation between two real (non-generated) images. Additionally, we evaluate reconstruction quality using two standard metrics: LPIPS Zhang et al. (2018) and PSNR. Figure 4(b) shows inversion-reconstruction consistency. our results confirm that the proposed approach preserves information with high quality.
@@ 3.2 Modular Style Transfer
Style transfer has been widely studied since Gatys et al. (2016) proposed optimizing image pixels to match style and content statistics. Johnson et al. (2016) introduced perceptual-loss training of feed-forward networks, making style transfer practical.
@@
We fix $g_{x}$ and $g_{y}$ and associate each style with a matrix $A_{\text{style}}$. Analogously to $A_{t}$, the matrix $A_{\text{style}}$ is produced by a hypernetwork that takes a style index as input (rather than time, as in the one-step setting). Then
@@
with results shown in Figure 5. In the first row, we interpolate between the mosaic and candy styles; in the second row between rain-princess and udine; and in the third row between candy and rain-princess. See App.E for additional transfers.

### v2

@@ Abstract
Neural networks are famously nonlinear. However, linearity is defined relative to a pair of vector spaces, $f:\mathcal{X}\to\mathcal{Y}$. Leveraging the algebraic concept of transport of structure, we propose a method to explicitly identify non-standard vector spaces where a neural network acts as a linear operator. When sandwiching a linear operator $A$ between two invertible neural networks, $f(x)=g_{y}^{-1}(Ag_{x}(x))$, the corresponding vector spaces $\mathcal{X}$ and $\mathcal{Y}$ are induced by newly defined addition and scaling actions derived from $g_{x}$ and $g_{y}$. We term this kind of architecture a Linearizer. This framework makes the entire arsenal of linear algebra, including SVD, pseudo-inverse, orthogonal projection and more, applicable to nonlinear mappings. Furthermore, we show that the composition of two Linearizers that share a neural network is also a Linearizer. We leverage this property and demonstrate that training diffusion models using our architecture makes the hundreds of sampling steps collapse into a single step. We further utilize our framework to enforce idempotency (i.e. $f(f(x))=f(x)$) on networks leading to a globally projective generative model and to demonstrate modular style transfer.
Deep Learning, Linearity, Invertible Networks, Diffusion
Nimrod Berman* Assaf Hallak* Assaf Shocher*†
Ben-Gurion University NVIDIA Technion
bermann@post.bgu.ac.il ahallak@nvidia.com assaf.sh@technion.ac.il
11footnotetext: Equal contribution.22footnotetext: A.S. is a Chaya Fellow, supported by the Chaya Career Advancement Chair.footnotetext: Code available at https://github.com/assafshocher/Linearizer
@@ 1 Introduction
In contrast, non-linear systems, while more expressive, often defy such a structure. Iterating nonlinear mappings can quickly lead to intractable dynamics; inversion may be ill-posed or undefined; and even simple compositional questions lack closed-form answers (Strogatz, 2024). Neural networks, the dominant modeling tool in modern machine learning, are famously nonlinear, placing their analysis and manipulation outside the reach of classical linear algebra (Hornik et al., 1989). As a result, tasks that are trivial in the linear setting, such as projecting onto a subspace, enforcing idempotency, or collapsing iterative procedures, become major challenges in the nonlinear regime that require engineered loss functions and optimization schemes. This motivates our approach: applying the principle of conjugation (change of coordinates) to deep neural networks to induce a latent space where the mapping becomes linear. If so, we gain access to linear methods without sacrificing nonlinear expressiveness.
@@ 2 Linearizer Framework
Linearity is not an absolute concept; it is a property defined relative to a pair of vector spaces, $f:\mathcal{X}\to\mathcal{Y}$. This motivates the use of transport of structure (Mac Lane and Birkhoff, 1999) to identify a pair of spaces $\mathcal{X}$ and $\mathcal{Y}$, for which a function that is nonlinear over standard Euclidean space is, in fact, perfectly linear. Recall that a vector space is comprised of a set of vectors (e.g. $\mathbb{R}^{N}$), a field of scalars (e.g., $\mathbb{R}$), and two fundamental operations: vector addition (++) and scalar multiplication ($\cdot$). We propose inducing new learnable vector spaces by redefining operations, while keeping vectors and field unchanged.
Figure 1: Left. The Linearizer structure (top) is a linear operation sandwiched between two invertible functions (bottom). Right. Vector addition and scalar multiplication define induced vector spaces for which $f$ is linear.
@@ 2.1 Definitions
We introduce a formalism we term Linearizer, in which the relevant vector spaces are immediately identifiable by construction as they are isomorphic to the Euclidean space. The architecture we propose can be trained from scratch or by distillation from an existing model. Our approach is made practical by invertible neural networks (Rezende and Mohamed, 2015; Dinh et al., 2015, 2017). We build our model by wrapping a linear operator, a matrix $A$, between two such invertible networks, $g_{x}$ and $g_{y}$:
Definition 2.1 (Linearizer).
@@
Definition 2.2 (Induced Vector Space Operations).
Let $g:V\to V$ be an invertible function. We define a new set of operations, according to the transport of structure, $\oplus$ and $\odot$, for any two vectors $v_{1},v_{2}\in V$ and any scalar $a\in\mathbb{R}$:
@@
$g_{x}$, $g_{y}$, and the core $A$ are learned jointly end-to-end. [^1] Specifically, in our implementation each of $g_{x}$ and $g_{y}$ uses 6 invertible blocks (Appendix G.1). The specific INN architecture – affine coupling layers with a tiny U-Net conditioner – is detailed in Appendix G.2 and Appendix G.4. The linear core $A$ is task-specific and its implementations are summarized in Appendix G.3.
@@ 2.2 Linearity
The input space $\mathcal{X}$ is defined by operations $(\oplus_{x},\odot_{x})$ induced by $g_{x}$, and the output space $\mathcal{Y}$ by $(\oplus_{y},\odot_{y})$ induced by $g_{y}$. This is a vector-space isomorphism which promises preservation of geometry.
Proposition 2.3.
$(V,\oplus_{g},\odot_{g})$ is a vector space over $\mathbb{R}$. This is a direct consequence of the isomorphism. However, we also provide full verification in Appendix A.
Proposition 2.4.
@@
Linearity is preserved in transport of structure, but we can also verify:
$f(a_{1}\odot_{x}x_{1}\oplus_{x}a_{2}\odot_{x}x_{2})$
$=g_{y}^{-1}(Ag_{x}(g_{x}^{-1}(a_{1}g_{x}(x_{1})+a_{2}g_{x}(x_{2}))))$
@@
$=a_{1}\odot_{y}f(x_{1})\oplus_{y}a_{2}\odot_{y}f(x_{2})$
@@ 2.3 Intuition
The Linearizer can be understood through several lenses.
Linear algebra analogy.
The Linearizer is analogous to an eigendecomposition or SVD. Just as a matrix is diagonal (i.e. simplest) in its eigenbasis, our function $f$ is a simple matrix multiplication in the “linear basis” defined by $g_{x}$ and $g_{y}$. This provides a natural coordinate system in which to analyze and manipulate the function. For example, repeated application of a Linearizer with a shared basis ($g_{x}=g_{y}=g$) is equivalent to taking a power of its core matrix:
$\begin{split}\mathbb{L}_{\{g,g,A\}}^{\circ N}(x)&=\underbrace{f(f(\cdots f}_{N\text{ times}}(x)))\\ =g^{-1}(A^{N}\cdot g(x))&=\mathbb{L}_{\{g,g,A^{N}\}}(x)\end{split}$ (4)
Geometric interpretation.
$g_{x}$ can be seen as a diffeomorphism. Equip the input with the pullback metric $g_{x}^{\ast}\langle\cdot,\cdot\rangle_{\mathbb{R}^{N}}$ induced by the Euclidean metric in latent space. With respect to this metric, straight lines in latent correspond to geodesics in input, and linear interpolation in the induced coordinates maps to smooth and semantically coherent curves in input space.
[…]
Figure 2: Comparison between multi-step and one-step flow matching. Panel labels: (a) multi-step Linear FM. (b) one-step Linear flow matching (FM).
@@ 2.4 Properties
We review how basic properties of linear transforms are expressed by Linearizers. All these properties follow directly from the vector space isomorphism (transport of structure). Yet, we explicitly derive them separately here.
@@ 2.4.1 Composition
$\begin{split}(f_{2}\circ f_{1})(x)&=g_{z}^{-1}(A_{2}g_{y}(g_{y}^{-1}(A_{1}g_{x}(x))))\\ &=g_{z}^{-1}((A_{2}A_{1})\cdot g_{x}(x))\end{split}$ (5)
@@ 2.4.2 Inner Product and Hilbert Space
Definition 2.6 (Induced Inner Product).
@@
Proposition 2.7 (Induced spaces are Hilbert).
Let $g:V\to\mathbb{R}^{n}$ be a smooth bijection and endow $V$ with the induced vector space operations $(\oplus_{g},\odot_{g})$ and inner product $\langle u,v\rangle_{g}\;:=\;\langle g(u),\,g(v)\rangle_{\mathbb{R}^{n}}.$ Then $(V,\langle\cdot,\cdot\rangle_{g})$ is a Hilbert space. This is a characteristic of the transport of structure. For explicit proof See Appendix C.
@@ 2.4.3 Transpose
$=\langle x,g_{x}^{-1}(A^{\top}g_{y}(y))\rangle_{g_{x}}=\langle x,f^{\top}(y)\rangle_{g_{x}}.\qed$ (9)
@@ 2.4.4 Singular Value Decomposition
Spectral properties are transferred in isomorphic change of coordinates. We show this explicitly in Appendix B.
3 Theoretical Analysis
Before demonstrating applications, we address two fundamental theoretical questions: the expressive capacity of the Linearizer and the functional necessity of the core matrix.
3.1 Expressiveness
A natural concern is whether enforcing a linear bottleneck restricts the model’s capacity. Strictly speaking, the architecture does impose topological constraints on the global function $f$. For example, the kernel (null space) of $f$ is determined entirely by the kernel of $A$. Since linear subspaces are either trivial or infinite-dimensional, a Linearizer cannot, for instance, map exactly three distinct points to zero while mapping a fourth point to non-zero, as this would violate the subspace properties of the kernel in the latent space.
Despite this topological constraint, we prove that the Linearizer is expressive enough to fit any finite training set. In Appendix H.5 (Theorem H.4), we formally prove that for any finite dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, there exist invertible networks $g_{x},g_{y}$ and a linear map $A$ such that the Linearizer satisfies $f(x_{i})=y_{i}$ to arbitrary precision.
How can the model be topologically constrained yet fit any dataset? The resolution lies in the degrees of freedom available outside the data support. For example, the required infinite null-space can lie outside of the data. This logic also applies to infinite continuous manifolds of lower dimension than the space, allowing generalization on high-dimensional data, such as images or text embeddings that often reside on a low-dimensional manifold
3.2 On the Role of the Core Operator A
Given the power of $g$, what is the role of the matrix $A$? We analyze this in two regimes.
The General Regime ($g_{x}\neq g_{y}$). Let the SVD of $A$ be $U\Sigma V^{T}$. In the general case where $g_{x}$ and $g_{y}$ are distinct, we could define new invertible networks $g^{\prime}_{y}=g_{y}\circ U$ and $g^{\prime}_{x}=V^{T}\circ g_{x}$. The function would then be $f(x)=(g^{\prime}_{y})^{-1}(\Sigma g^{\prime}_{x}(x))$. If we further allow the networks to absorb the scaling defined by the singular values in $\Sigma$, then the core operator could be reduced to a diagonal matrix of zeros and ones (note that this does not imply idempotency unless $g_{x}=g_{y}$). In this sense, the fundamental role of $A$ is simply to define the rank of the function $f$.
The Constrained Regime ($g_{x}=g_{y}=g$). In applications like Flow Matching or Idempotency, the input and output spaces are identical so $g_{x}=g_{y}$. Here, the two networks are not independent. To preserve the structure $f(x)=g^{-1}(Ag(x))$, any transformation absorbed by $g$ must be inverted by $g^{-1}$. This rules out $g$ absorbing SVD elements $U,V$ as they are not necessarily mutual inverses. For diagonalizable matrices however, eigen-decomposition can be applied. Then, $A=Q\Lambda Q^{-1}$, where one could define $g^{\prime}=Q^{-1}\circ g$ and $f(x)=(g^{\prime})^{-1}(\Lambda g^{\prime}(x))$. Here, $A$ determines the function’s eigenvalues—a stronger role than just its rank—but this is
Practical Necessity: Functional Diversity. Does the above analysis suggest that we can always use a diagonal $A$? Crucially, this analysis only applies to learning a single static function. However, our applications typically utilize ”families” of Linearizers, that share the same $gx,gy$ with different $A$ matrices. For example our flow-matching application employs a different A for each timestep. The expressiveness inside such family is dictated by the expressiveness of $A$ which would be significantly reduced if it were diagonal.
@@ 4 Applications
[…]
Figure 3: One-step inversion examples: Left and right (in red): original (not generated) data $x_{1}$ and $x_{2}$. Intermediate images obtained by latent interpolation. See Appendix E.2 for more and higher-resolution results.
@@ 4.1 One-Step Flow Matching
Flow matching (Lipman et al., 2023; Song et al., 2021b) trains a network to predict the velocity field that transports noise samples to data points. In our setting, diffusion models (Ho et al., 2020) and flow matching (Lipman et al., 2023) are, from an algorithmic standpoint, equivalent under standard assumptions (Gao et al., 2024); hence we use the two frameworks interchangeably. Traditionally, this velocity must be integrated over time with many small steps, making generation slow. We train flow matching using our Linearizer as the backbone, with a single vector space ($g_{x}=g_{y}=g$). The key property that helps us achieve one-step generation is the closure of linear operators. Applying the model sequentially with some simple linear operations in between steps is a chain of linear operations. This chain can be collapsed to a single linear operator so that the entire trajectory is realized in a single step.
@@
$\begin{split}\mathcal{L}&=\mathbb{E}_{x_{0},x_{1},t}\,\bigl\|\,v\ominus f(x_{t},t)\,\bigr\|^{2}\\ &=\mathbb{E}_{x_{0},x_{1},t}\,\bigl\|\,g^{-1}\Bigl(g(x_{1})-g(x_{0})-A_{t}g(x_{t})\Bigr)\,\bigr\|^{2}\end{split}$ (13)
@@
Thus, what was originally a long sequence of updates collapses into a single multiplication. In practice, higher-order integration such as Runge–Kutta (Runge, 1895; Kutta, 1901) may be used to calculate $B$ more accurately (see Appendix F). $B$ is calculated only once, after training. Then generation requires only a single feedforward activation of our model on noise, regardless of the number of steps discretizing the ODE.
@@
The operator $A$ is produced by an MLP whose input is $t$. To avoid huge matrix multiplications we build $A$ as a low-rank matrix, by multiplying two rectangular matrices. $g$ does not take $t$ as input. Full implementation details are provided in App. G.
@@
Inversion and interpolation.
A fundamental limitation of flow models is that, in contrast to VAEs (Kingma and Welling, 2014), they lack a natural encoder: they cannot map data back into the prior (noise) space, and thus act only as decoders. As a result, inversion methods for diffusion (Dhariwal and Nichol, 2021; Song et al., 2021a; Huberman-Spiegelglas et al., 2024) have become an active research area. However, these techniques are approximate and often suffer from reconstruction errors, nonuniqueness, or computational overhead. Our framework connects diffusion-based models with encoding ability.
Lemma 4.1 (Moore–Penrose pseudoinverse of a Linearizer).
$f^{\dagger}(y)=g_{x}^{-1}(A^{\dagger}g_{y}(y))$ (18)
Proof.
We verify the four Penrose equations (Penrose, 1955) in Appendix D. ∎
This property enables the exact encoding of data in the latent space. For example, given two data points $x_{1},x_{2}$, we can encode them via $z^{a}=(1-a)\,f^{\dagger}(x_{1})\;+\;a\,f^{\dagger}(x_{2}),$ and decode back by $\hat{x}^{a}=f(z^{a}).$ Figure 3 shows latent interpolation between two real (non-generated) images. Additionally, we evaluate reconstruction quality using two standard metrics: LPIPS (Zhang et al., 2018) and PSNR. Figure 4(b) shows inversion-reconstruction consistency. Our results confirm high-quality information preservation.
@@ 4.2 Modular Style Transfer
Style transfer has been widely studied since (Gatys et al., 2016) proposed optimizing image pixels to match style and content statistics. (Johnson et al., 2016) introduced perceptual-loss training of feed-forward networks, making style transfer practical.
@@
We fix $g_{x}$ and $g_{y}$ and associate each style with a matrix $A_{\text{style}}$. The matrix $A_{\text{style}}$ is produced by a hypernetwork that takes a style index as input (rather than time, as in the one-step setting). Then
@@
with results shown in Figure 5. Rows 1–3 interpolate between mosaic/candy, rain-princess/udnie, and candy/rain-princess, respectively. See App. E for additional transfers.